# Personalized Spiking Neural Networks with Ferroelectric Synapses for EEG Signal Processing

### v2

@@ 4 Conclusion
From a BCI perspective, the ability to perform subject-specific adaptation directly on hardware is particularly significant, as it addresses the inherent non-stationarity and inter-subject variability of EEG signals. Rather than relying on fully subject-agnostic models or repeated offline retraining, programmable memristive synapses enable post-deployment personalization with limited energy and endurance overhead. This capability is critical for wearable and implantable BCI systems intended for long-term, real-world operation. While this study focuses on motor imagery decoding, the proposed deployment strategies and learning mechanisms are broadly applicable to other neural and wearable bio-signal processing tasks and should be further investigated in future work. Future efforts will focus on scaling to larger networks and full on-hardware deployment, further bridging the gap between brain-inspired algorithms and practical, personalized neuromorphic hardware systems.
@@
The PhysioNet EEG dataset used in this study is publicly available at https://physionet.org/content/
eegmmidb/1.0.0/.

### v3

@@ 4 Conclusion
From a BCI perspective, the ability to perform subject-specific adaptation directly on hardware is particularly significant, as it addresses the inherent non-stationarity and inter-subject variability of EEG signals. Rather than relying on fully subject-agnostic models or repeated offline retraining, programmable memristive synapses enable post-deployment personalization with limited energy and endurance overhead. This capability is critical for wearable and implantable BCI systems intended for long-term, real-world operation. The deployment strategies and learning mechanisms developed here are not limited to motor-imagery decoding, but are broadly relevant to other neural and wearable bio-signal processing tasks. Future work will therefore focus on scaling to larger networks and full on-hardware implementations, further bridging the gap between brain-inspired algorithms and practical, personalized neuromorphic hardware systems.
@@
The PhysioNet EEG dataset used in this study is publicly available at:
https://physionet.org/content/eegmmidb/1.0.0/.
The experimental data displayed in Fig. 2 and Fig. 3 is publicly available at:
https://doi.org/10.5281/zenodo.19928453